OpenStudy (anonymous):

Jane makes party bags for her upcoming birthday party. She puts 1 1/2 Hersheys kisses, 2 1/4 cups of starbursts and a 1/3 cup of peppermints. If she plans to put 1/2 cup in each bag, how many complete bags can she make?

OpenStudy (ghazi):

8

OpenStudy (anonymous):

Explain,will u plz?

OpenStudy (ghazi):

we have done it yesterday , add all the components, take the LCM and divide by 2 because 1 cup comes in 2 bags

OpenStudy (ghazi):

\[1\frac{ 1 }{ 2 }=\frac{ 1*2+1 }{ 2 }=\frac{ 3 }{ 2 }\] similarly , 2 1/4= 9/4 and 1/3 cups of peppermint so \[\frac{ 3 }{ 2 }+\frac{ 9 }{ 4 }+\frac{ 1 }{ 3 }=\frac{ 6*3+9*3+4*1 }{ 12 }=\frac{ 49 }{ 12 }\] now 1/2 cup is stored in 1 bag so 1 cup will be stored in 1/1/2=2 bags therefore 49/12 cups will be in \[\frac{ 49 }{ 12 }*2=\frac{ 49 }{ 6 }=8 \]

OpenStudy (anonymous):

let me read. 1 sec.

OpenStudy (anonymous):

Mk, I get it.:)

OpenStudy (anonymous):

Thank u

OpenStudy (ghazi):

:D YW

OpenStudy (anonymous):

:))))

OpenStudy (ghazi):

dont laugh teeth would fall out , old lady

OpenStudy (anonymous):

O yea. One did just now :D

OpenStudy (ghazi):

thats bad :)

OpenStudy (anonymous):

Yea. Wont be able to eat. *horrified*

OpenStudy (ghazi):

hmm that is a big trouble lol

OpenStudy (anonymous):

yea. must.stop.laughing!

OpenStudy (ghazi):

okay :)

OpenStudy (anonymous):

small smile --> all but one tooth fall out

OpenStudy (anonymous):

oops. gather the others, will u,lad? lol

OpenStudy (ghazi):

no old lady i wont lol

OpenStudy (anonymous):

:0 shomone hwelp!

OpenStudy (ghazi):

lol :) get your teeth i see ya later oldie

OpenStudy (anonymous):

alrighty:)